OpenStudy (anonymous):

Solve the system of equations. 5x + y = 9 3x + 2y = 4

OpenStudy (anonymous):

@iGreen @sammixboo

OpenStudy (anonymous):

ok imma use the substitution method starting with 5x + y = 9 5x + y = 9 5x + y - 5x = 9 - 5x y = 9 - 5x so y = 9 - 5x plug that in to this equation 3x + 2y = 4 3x + 2(9 - 5x) = 4 3x + 18 - 10x = 4 -7x + 18 = 4 -7x + 18 - 18 = 4 - 18 -7x = -14 -7x/-7 = -14/-7 x = 2 ok so plug in 2 where x is in 5x + y = 9 5*2 + y = 9 10 + y = 9 10 + y - 10 = 9 - 10 y = -1 so the solution for the two equations is (2, -1) right?

OpenStudy (anonymous):

i am going to check 5x + y = 9 5*2 + -1 = 9 10 + -1 = 9 9 = 9 3*2 + 2*-1 = 4 6 + -2 = 4 4 = 4 @iGreen @sammixboo

OpenStudy (anonymous):

i am guessing OS stopped loading for you @sammixboo

sammixboo (sammixboo):

Back! And yeah, OS just stopped. You are correct!

OpenStudy (igreen):

\(\Huge\color{lime} \checkmark\) Excellent work.